ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

19. History of Mathematics

1991 MS Classification: 1, 51, 52, 65, 76 Sastry, Seemanapalli V., Federal University of Paraiba, Brazil History of Mathematics: Yesterday, Today and Tomorrow Mathematics and its glory have been traced since the ages of Babilonia upto modern times. The Egyptian knowledge of Algebra seems to be much more limited than that of Babylonia. Among Hindu algebraists, Brahmagupta’s work of 630 B. C. goes beyond Diophantus. Bhaskara’s Arithmetic and Algebra of AD 1150 are the outstanding, giving rules for dealing with negative quantities. Muslim mid-ninth century Mathematics was highlighted. Medievial and modern European developments in Mathematics were traced, till the 21st century. Reference: Stillwell, John. Mathematics and Its history. Springer-Verlag, New York, 1989, pp. 1-371